OpenStudy (anonymous):

if 75 g of liquid water (C=4.18 J/g-°C) in a calorimeter changes temperature from 25°C to 75°C, how much heat was transferred? use the equation C=q/mΔT

OpenStudy (aaronq):

Use the equation provided

OpenStudy (anonymous):

can you walk me through it maybe? thats why i asked on here..... @aaronq

OpenStudy (aaronq):

I can help you correct any mistakes you make if you attempt it

OpenStudy (anonymous):

well its gonna be a long time because i dont know how to do it. thats why im asking for help.

OpenStudy (aaronq):

It's rather simple, assuming you know what the variables are. Plug in the variables where they go and solve. I don't mind helping you, hence why i'm here, but you have to put in some effort.

OpenStudy (anonymous):

assuming i know the variables. i dont. im not asking you to do it for me. but if you could idk start me off or something it would help alot

OpenStudy (aaronq):

Sure thing, the variables are the following: q = heat m= mass C = specific heat capacity (given in the question) \(\sf \Delta T\) is the change in temperature, explicitly, \(\sf \Delta T=T_{final}-T_{initial}\)

OpenStudy (anonymous):

the capacity is gonna be what 75? right?

OpenStudy (aaronq):

no, the specific heat capacity for water is: C=4.18 J/g-°C

OpenStudy (aaronq):

75 g is the mass

OpenStudy (anonymous):

i meant the other 75. but okay

OpenStudy (aaronq):

That's the temperature

OpenStudy (anonymous):

okay. writing out the equation

OpenStudy (anonymous):

i dont know

OpenStudy (aaronq):

lol you're so close. \(q= m*C*\Delta T\) Just plug the stuff in and solve, try it here

OpenStudy (anonymous):

i dont know what m or T is

OpenStudy (anonymous):

and that equation is different....

OpenStudy (aaronq):

m stands for mass, the mass of the water in this case \(\Delta T\) is the difference in temperature. Whats' the difference in temperature between 15 \(^oC\) and 25 \(^oC\)? \(\Delta T=25^oC-15^oC=10^oC\)

OpenStudy (aaronq):

The equation is the exact same, it's just rearranged with algebra

OpenStudy (anonymous):

well im not passing algebra soooo i was confused. sorry

OpenStudy (aaronq):

No need to apologize. I think algebra is necessary to solve these problems, a pre-requisite if you may. you're going to run into a lot of of these kinds of problems where you have to use algebra.

OpenStudy (anonymous):

15.7kJ